I have ack'ed all the alerts for now
|dbbackups: Update config for mysql backup monitoring||operations/puppet||production||+29 -19|
I believe to have fixed the issue, puppet was wrong after the package fix at T351491.
I am rerunning all snapshots to check it is now fixed. Logical backups will be reanalyzed to not lose their metadata. In theory backups continued running, just we lost temporarily its status and metadata.
So we should now wait and by the end of the day all alerts should go back to green.